# A Careful Examination of Large Language Model Performance on Grade School Arithmetic

### 01 May 2024 v1

@@ 1 Introduction
Improving reasoning in large language models (LLMs) is one of the most important directions of current research. As such, proper benchmarking of current LLM abilities is paramount for ensuring progress continues in the correct direction. Currently, the field typically relies on public benchmarks (e.g. GSM8k, MATH, MBPP, HumanEval, SWEBench Cobbe et al. (2021); Hendrycks et al. (2021b); Cobbe et al. (2021); Chen et al. (2021); Jimenez et al. (2024)). However, because LLMs are trained on large corpora of data scraped from the Internet, there are major concerns that such benchmarks may inadvertently include examples that closely resemble the questions found in such benchmarks. This contamination may result in models having weaker reasoning capabilities than otherwise believed, due to simply being able to repeat the correct answer that it previously encountered during pre- or post- training. To properly investigate the reasoning abilities of models, we commission GSM1k, a newly constructed collection of 1250 grade school level math problems designed to mirror that of GSM8k. We took extensive efforts to ensure that GSM1k had a similar distribution of difficulty to GSM8k to ensure an apples-to-apples comparison. These efforts are described in Section 3, alongside a detailed description of the data creation process. To mitigate worries about data contamination, we created GSM1k solely with human annotators, without assistance from any LLM or other synthetic data source.
@@ 2.1 Data Contamination
Because data contamination is a well known issue in the field Balloccu et al. (2024); Magar and Schwartz (2022); Sainz et al. (2023); Jacovi et al. (2023), model builders will frequently take great pains to minimize the likelihood of data contamination. For example, it is common to remove all data with too high of an n-gram overlap with the benchmark data Brown et al. (2020). Additionally, methods such as using embedding similarity attempt to remove all contaminated data that is too similar in embedding space to the dataset Shi et al. (2024). More recently, Srivastava et al. (2024) propose functional evaluations, where benchmarks are written in the form of functions that can generate an infinite number of specific evaluation datapoints, each with slightly different numbers. In this setup, whenever a language model is evaluated, functional evaluations generate a specific problem instance to evaluate the model on, which is then never used again. This reduces the worry of data contamination by ensuring that no datapoint is ever used twice. Like ours, their results indicate the LLMs may be severely overfit on benchmark data. The main advantage of our approach over a purely function based evaluation is that functional evaluations can only generate a tiny portion of the full problem space by producing variations of the same problem with slightly different numerical values.
@@ 3.2.1 Human Differentiation Rates
The first test we run is human distinguishability. We present human crowdworkers with a set of five questions, four of which were randomly selected from the original GSM8k dataset and one of which was selected from the newly created GSM1k dataset, and rewarded crowdworkers for finding the odd one out. In an audit conducted using 19 crowdworkers who were not involved in the problem creation process, we found that crowdworkers were able to correctly identify the lone GSM1k example 21.83% of the time out of 1205 attempts (20% is pure chance). Separately, we also tested several paper authors who had not yet seen the data and they were also unable to perform much better than random. This suggests minimal differences between GSM8k and GSM1k, at least as measured by the human eye.
@@ 3.2.2 Human Solve Rates
To ensure similar solve rates, we also asked crowdworkers to solve questions under time pressure. 14 crowdworkers who had not participated in the problem creation process attempted to solve as many GSM8k problems as they could in 15 minutes and rewarded them for the number of correct solutions. We repeated this exact setup for GSM1k. Crowdworkers were able to solve an average of $4.07\pm 0.93$ problems on the GSM8k dataset. They were able to solve $4.36\pm 1.11$ problems on the GSM1k dataset, where the error rates are the standard deviations of the evaluation. This suggests that GSM1k is comparable in difficulty (and perhaps even slightly easier) than GSM8k. As such, substantial decreases in model accuracy on GSM1k compared to GSM8k are likely not explainable due to differences in dataset difficulty.
@@ 4 Results
We select models to evaluate based on popularity. Additionally, we evaluated several lesser known models that sit near the top of the OpenLLMLeaderboard and discover evidence of Goodhart’s law: many of these models perform substantially worse on GSM1k, suggesting that they are primarily gaming the GSM8k benchmark rather than improving model reasoning capabilities. The full set of results, including the performance table for all models, can be found in Appendix D.
@@ 5.4 Lesson 4: Data Contamination Is Likely Not The Full Explanation for Overfitting
Nevertheless, data contamination is likely not the full story. We observe this via the presence of several outliers, which cause the $r^{2}=0.32$ value to be relatively low. Examining these outliers carefully reveals that the model with the lowest per-character log-likelihood (Mixtral-8x22b) and the model with the highest per-character log-likelihood (Mixtral-8x22b-Instruct) are not only variations of the same model, but also have similar levels of overfit Jiang et al. (2024). Perhaps more intriguingly, the most overfit model we discovered (Math-Shepherd-Mistral-7B-RL Yu et al. (2023)) had a relatively low per-character log-likelihood. Math Shepherd trains a reward model on process level data using synthetic data. As such, we hypothesize that the reward modelling process may have leaked information about the correct reasoning chains for GSM8k even if the problems themselves did not ever appear in the dataset. Finally, we observe that the Llema models Azerbayev et al. (2024) have both high log-likelihoods and minimal overfit. Since these models are open-source and their training data is known, it is known that several GSM8k problem instances appear in the training corpus, as described by the authors in their paper. Nevertheless, the authors find (and our study supports) that these few instances do not lead to serious overfitting. The existence of these outliers suggests that overfitting on GSM8k is not purely due to data contamination, but rather may be through other indirect means, such as model builders collecting data similar in nature to benchmarks as training data or selecting final model checkpoints based on performance on benchmarks, even if the model itself may have not seen the GSM8k dataset at any point via training. Conversely, the reverse is also true: small amounts of data contamination do not necessarily lead to overfitting.
@@ 7 Acknowledgements
We would like to thank Dan Hendrycks, Adi Ganesh, Akilesh Praveen, Andrea Jaba, Charlotte Zhuang, Will Zhou and Celia Chen for their helpful comments and suggestions.
@@ Appendix A Annotator Instructions
We provide the annotator instructions given to crowdworkers below
@@ Appendix B N-shot Prompt (examples selected randomly from GSM8k train)
84
Question: The highest temperature ever recorded on Earth was 136 degrees Fahrenheit and
the coldest temperature ever measured was -126 degrees Fahrenheit. If the average
temperature of Earth is 59, what would be the difference between the average temperature
on Earth and the average given the two extremes?
Answer:
@@ Appendix C Results with An Alternative Prompt
We report our results in Table 1. On average, we find that this prompt causes the gap between GSM8k and GSM1k to decrease by about 1% across all models. However, for some models (e.g. Mixtral-8x22B-v0), this reduces the amount of observed overfitting dramatically. While the exact cause of this difference is impossible to know, especially without access to model details such as their training set, our hypothesis is that prompting a model with GSM8k is more likely to activate the “memorization” portion of a model than if it is prompted by non GSM8k grade school math problems.
@@ Appendix D Results Table
We report our full results in Table 1. Models are sorted by the difference in performance between GSM8k and GSM1k. Because all models are evaluated using the standard LM Evaluation Harness prompt and evaluation format, model performance on GSM8k may not match reported benchmark numbers. In particular, answers that do not match the 5-shot example format are marked incorrect even if they are otherwise “correct.” Our focus is primarily on the difference between GSM8k and GSM1k performance, holding evaluation setting constant. Alternative prompt results are also included. For details, see Appendix C.
Model Diff GSM8k GSM1k GSM8k_alt GSM1k_alt Diff_alt
math-shepherd-mistral-7b-rl 0.135 0.745 0.611 0.782 0.645 0.138
Xwin-Math-13B-V1 0.101 0.631 0.529 0.643 0.567 0.076
Xwin-Math-7B-V1 0.101 0.529 0.428 0.513 0.456 0.058
Mixtral-8x22B-Instruct-v0 0.099 0.859 0.760 0.885 0.843 0.042
Phi-3-mini-4k-instruct 0.098 0.782 0.684 0.801 0.763 0.039
deepseek-llm-67b-base 0.093 0.615 0.522 0.656 0.594 0.061
Mixtral-8x22B-v0 0.092 0.769 0.677 0.810 0.791 0.019
gemma-7b 0.077 0.541 0.464 0.544 0.518 0.026
phi-2 0.074 0.569 0.495 0.554 0.513 0.041
Yi-6B-Chat 0.073 0.408 0.334 0.423 0.362 0.061
mistral-small-latest 0.072 0.790 0.718 0.782 0.751 0.031
Yi-34B-Chat 0.071 0.641 0.569 0.647 0.584 0.062
Smaug-34B-v0 0.069 0.757 0.688 0.745 0.700 0.046
falcon-180B-chat 0.065 0.610 0.545 0.622 0.597 0.025
codegemma-7b 0.063 0.479 0.416 0.515 0.459 0.057
CodeLlama-70b-Python-hf 0.062 0.462 0.399 0.493 0.439 0.054
command 0.062 0.445 0.383 0.446 0.397 0.049
Meta-Llama-3-8B-Instruct 0.062 0.752 0.690 0.772 0.746 0.026
Smaug-2-72B 0.061 0.799 0.738 0.812 0.755 0.057
CodeLlama-70b-hf 0.059 0.486 0.427 0.500 0.460 0.039
Phi-3-mini-128k-instruct 0.058 0.742 0.683 0.817 0.752 0.065
CodeLlama-34b-Instruct-hf 0.058 0.415 0.357 0.400 0.392 0.007
Meta-Llama-3-8B 0.057 0.499 0.442 0.547 0.529 0.018
codegemma-7b-it 0.055 0.526 0.471 0.515 0.459 0.056
phi-1 0.048 0.324 0.276 0.324 0.276 0.048
Mixtral-8x7B-v0 0.048 0.579 0.530 0.607 0.601 0.006
Phind-CodeLlama-34B-v2 0.048 0.416 0.368 0.403 0.376 0.026
Mixtral-8x7B-Instruct-v0 0.046 0.641 0.594 0.686 0.657 0.029
gemma-7b-it 0.044 0.335 0.291 0.248 0.240 0.008
mistral-medium-latest 0.038 0.790 0.752 0.789 0.764 0.025
vicuna-33b-v1 0.038 0.379 0.341 0.441 0.417 0.023
CodeLlama-13b-Python-hf 0.030 0.214 0.183 0.221 0.194 0.027
OpenMath-Llama-2-70b-hf 0.029 0.171 0.142 0.165 0.126 0.039
Mistral-7B-v0 0.029 0.381 0.353 0.421 0.415 0.007
CodeLlama-34b-hf 0.028 0.349 0.322 0.324 0.305 0.019
CodeLlama-13b-Instruct-hf 0.027 0.277 0.250 0.287 0.276 0.012
CodeLlama-70b-Instruct-hf 0.025 0.497 0.471 0.519 0.477 0.041
Llama-2-7b-hf 0.023 0.141 0.118 0.142 0.127 0.015
Meta-Llama-3-70B 0.022 0.811 0.789 0.807 0.842 -0.035
Llama-2-70b-hf 0.021 0.537 0.516 0.572 0.566 0.006
Meta-Llama-3-70B-Instruct 0.020 0.896 0.876 0.899 0.885 0.015
gemini-1.5 0.018 0.897 0.879 0.908 0.892 0.016
pythia-12b 0.016 0.036 0.021 0.050 0.027 0.024
CodeLlama-13b-hf 0.015 0.227 0.212 0.218 0.215 0.003
CodeLlama-34b-Python-hf 0.013 0.306 0.293 0.340 0.281 0.059
gemma-2b 0.013 0.176 0.163 0.191 0.178 0.013
dbrx-base 0.013 0.727 0.715 0.718 0.685 0.033
deepseek-coder-33b-instruct 0.010 0.419 0.409 0.421 0.392 0.029
CodeLlama-7b-Python-hf 0.008 0.127 0.119 0.123 0.119 0.003
gpt-3.5 0.007 0.760 0.753 0.742 0.759 -0.016
gpt2-xl 0.004 0.021 0.017 0.022 0.018 0.004
gpt-neox-20b 0.003 0.059 0.056 0.076 0.070 0.005
gemini-pro 0.002 0.791 0.789 0.688 0.687 0.000
CodeLlama-7b-hf 0.002 0.127 0.125 0.116 0.095 0.021
CodeLlama-7b-Instruct-hf -0.000 0.186 0.186 0.187 0.164 0.023
mistral-large-latest -0.000 0.853 0.853 0.854 0.857 -0.003
gpt-4-turbo -0.000 0.898 0.898 0.847 0.849 -0.002
claude-3-haiku-20240307 -0.000 0.785 0.785 0.791 0.791 0.000
gemma-2b-it -0.001 0.111 0.112 0.099 0.106 -0.007
Llama-2-13b-hf -0.004 0.231 0.235 0.281 0.279 0.002
claude-2 -0.007 0.887 0.894 0.836 0.836 0.001
gpt-4 -0.007 0.910 0.918 0.919 0.904 0.014
deepseek-math-7b-rl -0.022 0.185 0.207 0.754 0.646 0.108
claude-3-opus-20240229 -0.023 0.802 0.825 0.830 0.857 -0.026
claude-3-sonnet-20240229 -0.024 0.720 0.744 0.713 0.724 -0.011

### 03 May 2024 v3

@@ 1 Introduction
Improving reasoning in large language models (LLMs) is one of the most important directions of current research. As such, proper benchmarking of current LLM abilities is paramount for ensuring progress continues in the correct direction. Currently, the field typically relies on public benchmarks such as GSM8k Cobbe et al. (2021), MATH Hendrycks et al. (2021b), MBPP Austin et al. (2021), HumanEval Chen et al. (2021), SWEBench Jimenez et al. (2024)). However, because LLMs are trained on large corpora of data scraped from the Internet, there are major concerns that such benchmarks may inadvertently include examples that closely resemble the questions found in such benchmarks. This contamination may result in models having weaker reasoning capabilities than otherwise believed, due to simply being able to repeat the correct answer that it previously encountered during pre- or post- training. To properly investigate the reasoning abilities of models, we commission GSM1k, a newly constructed collection of 1250 grade school level math problems designed to mirror that of GSM8k. We took extensive efforts to ensure that GSM1k had a similar distribution of difficulty to GSM8k to ensure an apples-to-apples comparison. These efforts are described in Section 3, alongside a detailed description of the data creation process. To mitigate worries about data contamination, we created GSM1k solely with human annotators, without assistance from any LLM or other synthetic data source.
@@ 2.1 Data Contamination
Because data contamination is a well known issue in the field Balloccu et al. (2024); Magar and Schwartz (2022); Sainz et al. (2023); Jacovi et al. (2023); Xu et al. (2024), model builders will frequently take great pains to minimize the likelihood of data contamination. For example, it is common to remove all data with too high of an n-gram overlap with the benchmark data Brown et al. (2020). Additionally, methods such as using embedding similarity attempt to remove all contaminated data that is too similar in embedding space to the dataset Shi et al. (2024).
Xu et al. (2024) propose using similar variants of a benchmark questions to detect if models favor the original wording as a proxy for data contamination. Srivastava et al. (2024) propose functional evaluations, where benchmarks are written in the form of functions that can generate an infinite number of specific evaluation datapoints, each with slightly different numbers. In this setup, whenever a language model is evaluated, functional evaluations generate a specific problem instance to evaluate the model on, which is then never used again. This reduces the worry of data contamination by ensuring that no datapoint is ever used twice. Like ours, their results indicate the LLMs may be severely overfit on benchmark data. The main advantage of our approach over a purely function based evaluation is that functional evaluations can only generate a tiny portion of the full problem space by producing variations of the same problem with slightly different numerical values. Their results also suggest substantial amounts of data contamination, including for frontier models, in the MATH dataset.
@@ 3.2.1 Human Differentiation Rates
The first test we run is human distinguishability. We present human annotators with a set of five questions, four of which were randomly selected from the original GSM8k dataset and one of which was selected from the newly created GSM1k dataset, and rewarded annotators for finding the odd one out. In an audit conducted using 19 annotators who were not involved in the problem creation process, we found that annotators were able to correctly identify the lone GSM1k example 21.83% of the time out of 1205 attempts (20% is pure chance). Separately, we also tested several paper authors who had not yet seen the data and they were also unable to perform much better than random. This suggests minimal differences between GSM8k and GSM1k, at least as measured by the human eye.
@@ 3.2.2 Human Solve Rates
To ensure similar solve rates, we also asked annotators to solve questions under time pressure. 14 annotators who had not participated in the problem creation process attempted to solve as many GSM8k problems as they could in 15 minutes and were rewarded based on the number of problems they solved. We repeated this exact setup for GSM1k. Annotators were able to solve an average of $4.07\pm 0.93$ problems on the GSM8k dataset. They were able to solve $4.36\pm 1.11$ problems on the GSM1k dataset, where the error rates are the standard deviations of the evaluation. This suggests that GSM1k is comparable in difficulty (and perhaps even slightly easier) than GSM8k. As such, substantial decreases in model accuracy on GSM1k compared to GSM8k are likely not explainable due to differences in dataset difficulty.
@@ 4 Results
We select models to evaluate based on popularity. Additionally, we evaluated several lesser known models that sit near the top of the OpenLLMLeaderboard and discover evidence of Goodhart’s law: many of these models perform substantially worse on GSM1k, suggesting that they are primarily gaming the GSM8k benchmark rather than improving model reasoning capabilities. The full set of results, including the performance table for all models, can be found in Appendix D. For fair comparison, we partition the models by performance on GSM8k and compare them to other models which perform similarly (Figures 5, 6, 7).
@@ 5.4 Lesson 4: Data Contamination Is Likely Not The Full Explanation for Overfitting
Nevertheless, data contamination is likely not the full story. We observe this via the presence of several outliers, which cause the $r^{2}=0.32$ value to be relatively low. Examining these outliers carefully reveals that the model with the lowest per-character log-likelihood (Mixtral-8x22b) and the model with the highest per-character log-likelihood (Mixtral-8x22b-Instruct) are not only variations of the same model, but also have similar levels of overfit Jiang et al. (2024). Perhaps more intriguingly, the most overfit model we discovered (Math-Shepherd-Mistral-7B-RL Yu et al. (2023)) had a relatively low per-character log-likelihood. Math Shepherd trains a reward model on process level data using synthetic data. As such, we hypothesize that the reward modelling process may have leaked information about the correct reasoning chains for GSM8k even if the problems themselves did not ever appear in the dataset. Finally, we observe that the Llema models Azerbayev et al. (2024) have both high log-likelihoods and minimal overfit. These models are open-sourced alongside their training data, and the authors report finding a very small number of GSM8k examples in the training corpus. Nevertheless, they also find (and our study supports) that these few instances do not lead to overfitting. The existence of these outliers suggests that overfitting on GSM8k is not purely due to data contamination, but rather may be through other indirect means, such as model builders collecting data similar in nature to benchmarks as training data or selecting final model checkpoints based on performance on benchmarks, even if the model itself may have not seen the GSM8k dataset at any point via training. Conversely, the reverse is also true: small amounts of data contamination do not necessarily lead to overfitting.
@@ 7 Acknowledgements
We would like to thank Dan Hendrycks, Adi Ganesh, Akilesh Praveen, Andrea Jaba, Charlotte Zhuang, Will Zhou, Celia Chen and Kamilė Lukošiūtė for their helpful comments and suggestions.
@@ Appendix A Annotator Instructions
We provide the annotator instructions given below.
@@ Appendix B N-shot Prompt (examples selected randomly from GSM8k train)
#### 84
@@ Appendix C Results with An Alternative Prompt
We report our results in Table 1. On average, we find that this prompt causes the gap between GSM8k and GSM1k to decrease by about 1% across all models. However, for some models (e.g. Mixtral-8x22B-v0.1), this reduces the amount of observed overfitting dramatically. While the exact cause of this difference is impossible to know, especially without access to model details such as their training set, our hypothesis is that prompting a model with GSM8k is more likely to activate the “memorization” portion of a model than if it is prompted by non-GSM8k grade school math problems.
@@ Appendix D Results Table
We report our full results in Table 1. Models are sorted by the difference in performance between GSM8k and GSM1k. Because all models are evaluated using the standard LM Evaluation Harness prompt and evaluation format, model performance on GSM8k may not match reported benchmark numbers. In particular, answers that do not match the 5-shot example format are marked incorrect even if they are otherwise “correct.” Our focus is primarily on the difference between GSM8k and GSM1k performance, holding evaluation setting constant. The Z-score and p-value are calculated for a two-tailed two proportion Z-test. Alternative prompt results are also included. For details, see Appendix C.
Table 1: Standard Prompt
Model Diff GSM8k GSM1k Z-score p-value
math-shepherd-mistral-7b-rl 0.134 0.745 0.611 7.322 0.000
Xwin-Math-13B-V1.0 0.101 0.631 0.529 5.196 0.000
Xwin-Math-7B-V1.0 0.101 0.529 0.428 5.131 0.000
Mixtral-8x22B-Instruct-v0.1 0.099 0.859 0.760 6.403 0.000
Phi-3-mini-4k-instruct 0.098 0.782 0.684 5.600 0.000
deepseek-llm-67b-base 0.093 0.615 0.522 4.771 0.000
Mixtral-8x22B-v0.1 0.092 0.770 0.677 5.257 0.000
gemma-7b 0.077 0.541 0.464 3.879 0.000
phi-2 0.074 0.569 0.495 3.766 0.000
Yi-6B-Chat 0.073 0.408 0.334 3.851 0.000
mistral-small-latest 0.072 0.790 0.718 4.218 0.000
Yi-34B-Chat 0.071 0.641 0.569 3.683 0.000
Smaug-34B-v0.1 0.069 0.757 0.688 3.931 0.000
gemma-1.1-7b-it 0.069 0.522 0.453 3.485 0.000
falcon-180B-chat 0.065 0.610 0.545 3.320 0.001
codegemma-7b 0.063 0.479 0.416 3.217 0.001
CodeLlama-70b-Python-hf 0.063 0.462 0.399 3.198 0.001
command 0.062 0.445 0.383 3.221 0.001
Meta-Llama-3-8B-Instruct 0.062 0.752 0.690 3.532 0.000
Smaug-2-72B 0.061 0.799 0.738 3.697 0.000
CodeLlama-70b-hf 0.059 0.486 0.427 2.989 0.003
Phi-3-mini-128k-instruct 0.058 0.741 0.683 3.264 0.001
CodeLlama-34b-Instruct-hf 0.058 0.415 0.357 3.013 0.003
Meta-Llama-3-8B 0.057 0.499 0.442 2.865 0.004
codegemma-7b-it 0.055 0.526 0.471 2.784 0.005
phi-1 0.048 0.324 0.276 2.683 0.007
Mixtral-8x7B-v0.1 0.048 0.578 0.530 2.450 0.014
Phind-CodeLlama-34B-v2 0.048 0.416 0.368 2.460 0.014
Mixtral-8x7B-Instruct-v0.1 0.046 0.641 0.594 2.411 0.016
gemma-7b-it 0.044 0.335 0.291 2.397 0.017
mistral-medium-latest 0.038 0.790 0.752 2.292 0.022
vicuna-33b-v1.3 0.038 0.379 0.341 2.020 0.043
CodeLlama-13b-Python-hf 0.030 0.214 0.183 1.942 0.052
OpenMath-Llama-2-70b-hf 0.029 0.171 0.142 2.072 0.038
Mistral-7B-v0.1 0.029 0.381 0.353 1.500 0.134
CodeLlama-34b-hf 0.028 0.350 0.322 1.497 0.134
CodeLlama-13b-Instruct-hf 0.027 0.277 0.250 1.560 0.119
Mistral-7B-Instruct-v0.2 0.027 0.426 0.399 1.383 0.167
CodeLlama-70b-Instruct-hf 0.025 0.497 0.471 1.287 0.198
Llama-2-7b-hf 0.023 0.141 0.118 1.766 0.077
Meta-Llama-3-70B 0.022 0.811 0.789 1.371 0.170
Llama-2-70b-hf 0.021 0.538 0.516 1.092 0.275
Meta-Llama-3-70B-Instruct 0.020 0.896 0.876 1.607 0.108
Mistral-7B-Instruct-v0.1 0.019 0.335 0.316 1.033 0.302
gemini-1.5-pro-preview-0409 0.018 0.897 0.879 1.423 0.155
pythia-12b 0.016 0.036 0.021 2.362 0.018
CodeLlama-13b-hf 0.015 0.227 0.212 0.899 0.369
CodeLlama-34b-Python-hf 0.013 0.306 0.293 0.746 0.456
gemma-2b 0.013 0.176 0.163 0.911 0.362
dbrx-base 0.013 0.727 0.715 0.715 0.474
deepseek-coder-33b-instruct 0.010 0.419 0.409 0.538 0.591
CodeLlama-7b-Python-hf 0.008 0.127 0.120 0.629 0.529
gpt-3.5-turbo 0.007 0.760 0.753 0.405 0.685
gpt2-xl 0.004 0.021 0.017 0.665 0.506
gpt-neox-20b 0.003 0.059 0.056 0.253 0.800
gemini-pro 0.002 0.792 0.789 0.119 0.905
CodeLlama-7b-hf 0.002 0.127 0.125 0.135 0.893
CodeLlama-7b-Instruct-hf -0.000 0.186 0.186 0.010 0.992
mistral-large-latest -0.000 0.853 0.853 0.009 0.993
gpt-4-turbo -0.000 0.898 0.898 0.004 0.997
claude-3-haiku-20240307 -0.000 0.785 0.785 -0.007 0.994
gemma-2b-it -0.001 0.111 0.112 -0.044 0.965
Llama-2-13b-hf -0.004 0.230 0.235 -0.283 0.777
claude-2.1 -0.007 0.887 0.894 -0.532 0.595
gpt-4 -0.007 0.911 0.918 -0.638 0.523
gemma-1.1-2b-it -0.008 0.104 0.112 -0.664 0.506
deepseek-math-7b-rl -0.022 0.185 0.207 -1.418 0.156
claude-3-opus-20240229 -0.023 0.802 0.825 -1.474 0.141
claude-3-sonnet-20240229 -0.024 0.719 0.744 -1.355 0.175
Table 2: Alternative Prompt
Model Diff GSM8k GSM1k Z-score p-value
math-shepherd-mistral-7b-rl 0.138 0.782 0.645 7.726 0.000
deepseek-math-7b-rl 0.108 0.754 0.646 5.979 0.000
Xwin-Math-13B-V1.0 0.076 0.643 0.567 3.966 0.000
Phi-3-mini-128k-instruct 0.065 0.817 0.752 4.029 0.000
gemma-1.1-7b-it 0.065 0.493 0.427 3.334 0.001
Yi-34B-Chat 0.062 0.647 0.584 3.266 0.001
deepseek-llm-67b-base 0.061 0.656 0.594 3.215 0.001
Yi-6B-Chat 0.061 0.423 0.362 3.187 0.001
CodeLlama-34b-Python-hf 0.059 0.340 0.281 3.216 0.001
Xwin-Math-7B-V1.0 0.058 0.513 0.456 2.903 0.004
codegemma-7b 0.057 0.516 0.459 2.896 0.004
Smaug-2-72B 0.057 0.812 0.755 3.497 0.000
codegemma-7b-it 0.056 0.516 0.459 2.855 0.004
CodeLlama-70b-Python-hf 0.054 0.493 0.439 2.722 0.006
command 0.049 0.446 0.396 2.513 0.012
phi-1 0.048 0.324 0.276 2.637 0.008
Smaug-34B-v0.1 0.046 0.745 0.700 2.607 0.009
Mixtral-8x22B-Instruct-v0.1 0.042 0.885 0.843 3.077 0.002
CodeLlama-70b-Instruct-hf 0.041 0.519 0.477 2.117 0.034
phi-2 0.041 0.554 0.513 2.103 0.035
CodeLlama-70b-hf 0.039 0.500 0.460 2.009 0.045
OpenMath-Llama-2-70b-hf 0.039 0.165 0.126 2.796 0.005
Phi-3-mini-4k-instruct 0.039 0.801 0.763 2.392 0.017
dbrx-base 0.033 0.718 0.685 1.837 0.066
Mistral-7B-Instruct-v0.2 0.033 0.436 0.403 1.680 0.093
mistral-small-latest 0.031 0.782 0.751 1.871 0.061
Mixtral-8x7B-Instruct-v0.1 0.029 0.686 0.657 1.540 0.124
deepseek-coder-33b-instruct 0.028 0.421 0.392 1.484 0.138
CodeLlama-13b-Python-hf 0.027 0.221 0.194 1.683 0.092
gemma-7b 0.026 0.544 0.518 1.318 0.188
Phind-CodeLlama-34B-v2 0.026 0.403 0.376 1.381 0.167
Meta-Llama-3-8B-Instruct 0.026 0.772 0.746 1.505 0.132
falcon-180B-chat 0.025 0.622 0.597 1.292 0.196
mistral-medium-latest 0.025 0.789 0.764 1.488 0.137
pythia-12b 0.023 0.050 0.027 3.111 0.002
vicuna-33b-v1.3 0.023 0.440 0.417 1.171 0.241
CodeLlama-7b-Instruct-hf 0.023 0.187 0.164 1.548 0.122
CodeLlama-7b-hf 0.021 0.116 0.095 1.781 0.075
Mixtral-8x22B-v0.1 0.019 0.810 0.791 1.173 0.241
CodeLlama-34b-hf 0.019 0.324 0.305 1.033 0.302
Meta-Llama-3-8B 0.018 0.547 0.529 0.904 0.366
gemini-1.5-pro-preview-0409 0.016 0.908 0.892 1.376 0.169
Llama-2-7b-hf 0.015 0.143 0.127 1.136 0.256
Meta-Llama-3-70B-Instruct 0.015 0.899 0.885 1.173 0.241
gpt-4 0.014 0.919 0.904 1.259 0.208
gemma-2b 0.013 0.191 0.178 0.879 0.380
CodeLlama-13b-Instruct-hf 0.012 0.287 0.276 0.684 0.494
gemma-7b-it 0.008 0.248 0.240 0.467 0.641
CodeLlama-34b-Instruct-hf 0.007 0.400 0.392 0.391 0.696
Mistral-7B-v0.1 0.007 0.422 0.415 0.325 0.745
Mixtral-8x7B-v0.1 0.006 0.607 0.601 0.336 0.737
Llama-2-70b-hf 0.006 0.572 0.566 0.309 0.757
gpt-neox-20b 0.005 0.076 0.070 0.527 0.598
gpt2-xl 0.004 0.022 0.018 0.645 0.519
CodeLlama-7b-Python-hf 0.003 0.123 0.120 0.281 0.779
CodeLlama-13b-hf 0.003 0.218 0.215 0.147 0.883
Llama-2-13b-hf 0.002 0.281 0.279 0.117 0.907
claude-2.1 0.001 0.836 0.836 0.016 0.987
gemini-pro 0.000 0.688 0.687 0.024 0.981
claude-3-haiku-20240307 0.000 0.792 0.791 0.019 0.985
gpt-4-turbo -0.002 0.847 0.849 -0.137 0.891
mistral-large-latest -0.003 0.854 0.857 -0.225 0.822
gemma-1.1-2b-it -0.006 0.084 0.090 -0.490 0.624
gemma-2b-it -0.007 0.099 0.106 -0.589 0.556
Mistral-7B-Instruct-v0.1 -0.007 0.347 0.354 -0.380 0.704
claude-3-sonnet-20240229 -0.011 0.713 0.724 -0.596 0.551
gpt-3.5-turbo -0.016 0.742 0.759 -0.946 0.344
claude-3-opus-20240229 -0.026 0.830 0.857 -1.855 0.064
Meta-Llama-3-70B -0.035 0.807 0.841 -2.322 0.020